Please help me solve this equation.
2(3n(n-2)) = 3(2n(n-1)+12

Answer by ikleyn(52777)


2(3n(n-2)) = 3(2n(n-1) + 12,

2*(3n^2 - 6n) = 3*(2n^2 -2n) + 12,

6n^2 - 12n = 6n^2 - 6n + 12.

-6n = 12.

n = -2.

Check:  Left side is 2*3*(-2)*(-4) = 48.   Right side is 3*2*(-2)*(-3) + 12 = 36 + 12 = 48.   Both sides are the same.

Answer.  The solution is  n = -2.